OpenStudy (anonymous):

evaluate the integral from negative infinity to infinity of x^3 dx

jimthompson5910 (jim_thompson5910):

hint: x^3 is an odd function

OpenStudy (anonymous):

@jim_thompson5910 that doesn't seem to relevant here. The integral doesn't converge (although the principal value integral does...).

jimthompson5910 (jim_thompson5910):

the symmetry is very relevant

jimthompson5910 (jim_thompson5910):

although, it seems like you'd have infinity - infinity, so I'm not sure

OpenStudy (anonymous):

Yeah the symmetry tells us that the PV integral converges to zero, but not the regular integral (which I'm assuming is what's being asked for).

OpenStudy (gorv):

it will be converted to even function after integral @SithsAndGiggles....@jim_thompson5910 is right

OpenStudy (gorv):

\[\int\limits_{-\infty}^{\infty}x^3*dx=[\frac{ x^4 }{ 4 }]\]

OpenStudy (gorv):

with given limit

OpenStudy (gorv):

\[\frac{ (\infty)^4 }{ 4 }-\frac{ (-\infty)^4 }{ 4 }\]

OpenStudy (gorv):

\[(-\infty)^4=(\infty)^4\]

OpenStudy (gorv):

so result will be zero

OpenStudy (anonymous):

@gorv Then I suppose \[\lim_{x\to\infty}(x^3-x^2)~\text{ is also }0~?\]

OpenStudy (anonymous):

Because as you say, \(\infty-\infty=0\).

OpenStudy (gorv):

well there is a difference between limit and integration

OpenStudy (gorv):

integration calculate area ..and both side having equal area

OpenStudy (gorv):

and for your kind info

OpenStudy (gorv):

\[\infty^4=\infty^4\]

OpenStudy (gorv):

\[\infty^3>\infty^2\]

OpenStudy (gorv):

well check again ..

OpenStudy (gorv):

\[\lim_{x \rightarrow \infty} (x^2-(-x)^2)\]

OpenStudy (gorv):

will be zero for sure ...

OpenStudy (gorv):

@SithsAndGiggles

OpenStudy (anonymous):

\[\large\int_{-\infty}^\infty x^3~dx=\lim_{a\to-\infty}\int_a^c x^3~dx+\lim_{b\to\infty}\int_c^b x^3~dx\] for some valid \(c\) in the domain of \(\large x^3\) (which would be any real number). This is how we treat improper integrals like this one - we must treat the limits of integration as limits themselves. Besides, integrals are infinite sums of infinitesimal areas, so exactly how is this different from a limit when the definition of the definite integral IS a limit? As for my counter-example (which apparently didn't land) - my point was that \(\infty-\infty\) is an indeterminate form. We can't say that one infinity necessarily has the same magnitude as another infinity, to put it in layman's terms. If my reasoning doesn't hold any merit, then perhaps Wolfram's calculation will convince you otherwise: http://www.wolframalpha.com/input/?i=Integrate%5Bx%5E3%2C%7Bx%2C-Infinity%2CInfinity%7D%5D